Two sided dice are tossed. Find the probability that the sum is 5 or less.
a. 1/3
b. 4/11
c. 5/12
d. 5/18

Answer by stanbon(75887)

Two sided dice are tossed. Find the probability that the sum is 5 or less.
a. 1/3
b. 4/11
c. 5/12
d. 5/18
-----------
Count the possible results that meet the condition:
(1,1),(1,2),(1,4)
(2,1),(2,2),(2,3)
(3,1),(3,2)
(4,1)
------------------
Probability(sum is 5 or less) = 9/36 = 1/4
==================
Cheers,
Stan H.